# Unraveling particle dark matter with Physics-Informed Neural Networks

### v1

@@ 1 Introduction
The Standard Model (SM) of particle physics, which describes the fundamental particles and their interactions, stands among the greatest achievements in theoretical physics. The predictions of the SM have been in remarkable agreement with experimental results. Namely, its last missing piece, the Higgs boson, was discovered in 2012, by ATLAS and CMS, at the Large Hadron Collider (LHC) [1]. Nonetheless, there are significant experimental evidences for new physics. One of the most striking examples of such new physics is the existence of dark matter (DM), established through indirect astrophysical measurements and cosmological observations (for a recent review see Ref. [2]). These indicate that about $27\%$ of the total matter in the Universe is DM, being about $5$ times more abundant than ordinary baryonic matter [3]. Precisely, the observed cold DM (CDM) relic abundance obtained by the Planck satellite data is $\Omega_{\text{CDM}}h^{2}=0.120\pm 0.001$ [3]. This data point assumes the $\Lambda$CDM model, which describes the evolution of our Universe, its matter content, structure formation and accelerated expansion following the Hubble law, based on the Big-Bang theory. A viable particle DM candidate must be electrically neutral, non-baryonic, stable and non-relativistic (cold). Since no such particle is found in the SM, we must explore beyond the SM extensions.
@@
The impact of ML has been so significant that the 2024 physics Nobel prize recognized the application of physics concepts to its foundations. In particular, deep learning techniques based on neural networks (NNs) have been very successful, solving a multitude of problems in different fields such as computer vision, natural language processing, game theory, science and engineering, among others. The application of various ML algorithms has also made its way into particle physics with an ever growing popularity [24]. Typically, these are incorporated into experiments, which generally require the analysis of large volumes of high-dimensional and highly-correlated data. But even from a theoretical standpoint, particle physics provides problems where ML has been shown to be useful – see e.g. Refs. [25, 26, 27, 28, 29]. In this work we will make use of an algorithm called Physics-Informed Neural Networks (PINNs) (for a review see Ref. [30]). Recent developments have sought to extend the performance of the basic PINN architecture, e.g. [31, 32, 33], by addressing the so-called failure modes of PINNs [34, 35]. The idea of the basic PINN is to make use of NNs as universal approximators [36], where prior knowledge of physics is fed into the learning process of the NNs. The aim is to solve partial and/or ordinary differential equations (PDEs/ODEs) [37]. The NN is able to achieve such task via the minimization of a loss function encoding the residuals of the given PDEs or ODEs, as well as initial and/or final and/or boundary conditions. Adding this information to the loss function optimizes the learning process of a classical NN. In fact, PINNs do not require large amounts of training data and can make use of automatic differentiation (AD) [38] to evaluate all differential operators entering such equations. PINNs have been used to tackle a multiplicity of problems, e.g. involving stiff equations [39], heat equations [40], equations of state of astrophysical objects [41, 42, 43], obtaining quasi-normal modes of black holes [44, 45, 46], and only very recently in the context of high-energy-physics [47, 48, 49]. One type of generic problem PINNs have been used for are the so-called forward problems. Namely, given a physical law, usually characterized by PDEs or ODEs, and initial/boundary conditions, a solution is found by the PINN, corresponding to a physical quantity. But there is another interesting class of problems that can be solved by PINNs: inverse problems. In this case, the observable quantities are given to the PINN which subsequently determines the physical laws and/or parameters of a given physical theory that can satisfy such observables. This is of particular interest, since PINNs can potentially be a tool to discover models and/or a given models’ parameter space that inherently satisfy experimental observables.
@@
Inspired by the above ideas, in this work, we solve the freeze-in DM BE in alternative cosmologies using PINNs. Our aim is to showcase how PINNs are a powerful technology for exploring theoretical physics models motivated by experimental data, by addressing:
@@ 2 Freeze-in dark matter in alternative cosmologies
with $S_{0}=2.89\times 10^{3}$ cm-3 being the entropy at present time, $\rho_{\text{crit}}^{0}=1.05h^{2}\times 10^{-5}$ GeV///cm3 the present critical energy density and $h=H_{0}/(100$ km///s///Mpc) the reduced Hubble constant [50]. The above needs to be confronted with the experimetally observed value $\Omega_{\text{CDM}}h^{2}=0.120\pm 0.001$ [3].
@@ 3 Recasting the Boltzmann equation for freeze-in dark matter
The performance of the PINN can be further improved if we normalize $W(z)$. Since the derivative $dY(x)/dx$ is always positive in Eq. (1) and the logarithm is a strictly monotonically increasing function, any initial condition (IC) we impose establishes the minimum of $W(z)$. Considering $z_{0}>0$ the point where such minimum occurs and $W_{0}\equiv W(z_{0})$, we rescale $W(z)$ as follows,
$W_{n}(z)=-W(z)/W_{0}\;.$ (14)
The minimum of this new function is $-1$, while its maximum is within the interval [1,0[[-1,0[[ - 1 , 0 [.
@@ 4 Forward problem: parametric PINN solution
where we fix $W_{0}=\ln 10^{-25}$. In order to improve the performance of PINNs in solving PDEs, some mechanisms to make loss weights adaptive have been put forward, such as learning rate annealing [34, 52] and soft attention [53]. Although the BE imposes similar challenges to those tackled in the aforementioned references, we found no advantages in implementing these mechanisms. Hence, $\lambda_{\text{BE}}$ and $\lambda_{\text{IC}}$, were kept constant throughout training. By manually adjusting these constants, we found better results when $\lambda_{\text{IC}}$ was fixed at a higher value than $\lambda_{\text{BE}}$, so we take $\lambda_{\text{BE}}=1$ and $\lambda_{\text{IC}}=10$.
@@
Figure 3: Parametric forward PINN solution: particle yield evolution $W_{n}(z)$ for Standard (top), RS (bottom left) and GB (bottom right) cosmology. Parametric dependence of $W_{n}(z)$ on $C$ is shown via a colormap. Dashed-red curves, obtained via FEM, indicate upper and lower limits of Standard cosmology results. Horizontal dotted-black line corresponds to $W_{n}$ value satisfying $\Omega h^{2}=0.12$ [3].
Figure 4: Relation between $W_{n}(z_{f})$ and $C$ for Standard (blue), GB (green) and RS (red) cosmology.
@@
In Fig. 3 we present the parametric PINN solution for the forward problem in Standard (top), RS (bottom left) and GB (bottom right) cosmology. The behavior of the yield $W_{n}(z)$ with respect to $C$ is represented by a colormap, while the red-dashed lines show the solution provided by solve_ivp, when $C=-14$ and $C=-9$, for Standard cosmology. We observe a nearly perfect agreement between the PINN and solve_ivp in the top plot, with the solution for $C\in[-14,-9]$ filling practically all the region between the red-dashed lines. This showcases the accuracy of the forward PINN in solving the BE. Moreover, we notice that for $C\sim-11.5$, in Standard cosmology, our FIMP DM candidate accounts for the observed DM relic abundance (horizontal dotted-black line). For GB cosmology one would need a decreased value of $C\sim-13.5$ (bottom right plot), while for RS an increased value $C>-9$ is required (bottom left). These results are consistent with the theoretical expectations outlined in Section 2 [see Eqs. (9) and (12)]. In addition, in Fig. 4, we present the yield value provided by the PINN for the last collocation point $W_{n}(z_{f})$ in terms of $C$, for the three cosmologies. The PINN confirms the expected linear relation between relic density and cross-section for FIMP DM [see Eqs. (9) and (12)]. Overall, compared to FEMs, the forward PINN provides a mesh-free method to solve the BE parametrically for cross-section values within an interval. Indeed, the NN is here a continuous function that approximates our target $W(z,C)$.
@@ 5.1 Discovering particle interaction cross sections
$\mathcal{L}_{\Omega h^{2}}=\left|W_{n}(z_{f})-W_{\Omega h^{2}}\right|\;,$ (23)
measures the absolute difference between the prediction of the PINN for the last collocation point $W_{n}(z_{f})$ and the observed DM relic density [see Eq. (13)]. Like $\lambda_{\text{BE}}$ and $\lambda_{\text{IC}}$, the weight $\lambda_{\Omega h^{2}}$ is an hyperparameter that should be tuned to optimize the PINN’s performance. As in the forward problem we have previously discussed, no improvements were found by making those loss weights adaptive. Manual adjustment leads us to set $\lambda_{\text{BE}}=1$ and $\lambda_{\text{IC}}=\lambda_{\Omega h^{2}}=10$.
@@
Figure 8: Inverse PINN solution: particle yield evolution $W_{n}(z)$ for Standard (top), RS (bottom left) and GB (bottom right) cosmology. Black-dashed curves were obtained via FEM.
In Figs. 7and 8, we present the PINN solution for the inverse problem of finding the FIMP DM particle yield evolution and determining the particle interaction strength for Standard (blue line), RS (red line) and GB (green line) cosmology. Namely, in Fig. 7, we show how $C$ evolves throughout training. The inverse PINN reaches the correct value for particle cross-sections after about $50-100$ thousand epochs for the three cosmologies being tested. Remarkably, Fig. 7 also shows that this can be achieved regardless of the initial $C$ being quite close (as in the case of the Standard cosmology) or far (as it happens in the alternative cosmologies) from the value that the PINN is supposed to discover. The PINN predictions for $C$ and corresponding cross-section [see Eq. (15)] are gathered in Table 1 for the three cosmologies. We also show the particle yield values for the last collocation point $W_{n}(z_{f})$ and relic density predictions stemming from our FEM BE solver when we provide it the $C$ values obtained by the PINN and their respective $\gamma$’s. This serves as a cross-check of the PINN results in obtaining the correct CDM relic abundance. In the last column of the table, the central value of the data is reproduced up to the fourth decimal place, showcasing the accuracy of the inverse PINN. Moreover, from Fig. 8, we also conclude that all cosmological scenarios reproduce the data $\Omega h^{2}=0.12$ (horizontal black-dashed line).
$\gamma$ $C_{\text{PINN}}$ $\langle\sigma v\rangle$ $\left(\text{GeV}^{-2}\right)$ $W_{n}(z_{f})_{\text{FEM}}$ $\Omega h^{2}_{\text{FEM}}$
@@
Table 1: Inverse PINN solution: finding $C$ [or equivalently $\langle\sigma v\rangle$ – see Eq. (15)] for GB ($\gamma=-2/3$), Standard ($\gamma=0$) and RS ($\gamma=2$) cosmology –see Figs. 7 and 8. Last two-columns indicate relic abundance values obtained via FEM.

### v2

@@ 1 Introduction
The Standard Model (SM) of particle physics, which describes the fundamental particles and their interactions, stands among the greatest achievements in theoretical physics. The predictions of the SM have been in remarkable agreement with experimental results. Namely, its last missing piece, the Higgs boson, was discovered in 2012, by ATLAS and CMS, at the Large Hadron Collider (LHC) [1]. Nonetheless, there are significant experimental evidences for new physics. One of the most striking examples of such new physics is the existence of dark matter (DM), established through indirect astrophysical measurements and cosmological observations (for a recent review see Ref. [2]). These indicate that about $27\%$ of the total energy in the Universe is DM, being about $5$ times more abundant than ordinary baryonic matter [3]. Precisely, the observed cold DM (CDM) relic abundance obtained by the Planck satellite data is $\Omega_{\text{CDM}}h^{2}=0.120\pm 0.001$ [3]. This data point assumes the $\Lambda$CDM model, which describes the evolution of our Universe, its matter content, structure formation and accelerated expansion following the Hubble law, based on the Big-Bang theory. A viable particle DM candidate must be electrically neutral, non-baryonic, stable and non-relativistic (cold). Since no such particle is found in the SM, we must explore beyond the SM extensions.
@@
The impact of ML has been so significant that the 2024 physics Nobel prize recognized the application of physics concepts to its foundations. In particular, deep learning techniques based on neural networks (NNs) have been very successful, solving a multitude of problems in different fields such as computer vision, natural language processing, game theory, science and engineering, among others. The application of various ML algorithms has also made its way into particle physics with an ever growing popularity [67]. Typically, these are incorporated into experiments, which generally require the analysis of large volumes of high-dimensional and highly-correlated data. But even from a theoretical standpoint, particle physics provides problems where ML has been shown to be useful – see e.g. Refs. [68, 69, 70, 71, 72, 73, 74, 75]. In this work we will make use of an algorithm called Physics-Informed Neural Networks (PINNs) (for a review see Ref. [76]). Recent developments have sought to extend the performance of the basic PINN architecture, e.g. [77, 78, 79], by addressing the so-called failure modes of PINNs [80, 81]. The idea of the basic PINN is to make use of NNs as universal approximators [82], where prior knowledge of physics is fed into the learning process of the NNs. The aim is to solve partial and/or ordinary differential equations (PDEs/ODEs) [83]. The NN is able to achieve such task via the minimization of a loss function encoding the residuals of the given PDEs or ODEs, as well as initial and/or final and/or boundary conditions. Adding this information to the loss function optimizes the learning process of a classical NN. In fact, PINNs do not require large amounts of training data and can make use of automatic differentiation (AD) [84] to evaluate all differential operators entering such equations. PINNs have been used to tackle a multiplicity of problems, e.g. involving stiff equations [85, 86, 87, 88, 79]; heat equations [89]; Navier-Stokes equations for fluid dynamics [90, 91]; modeling astrophysical objects with data and equations of state [92, 93, 94, 95, 96, 97, 98, 99]; studying neutrino flavor conversion in supernova [100, 101]; modeling DM halo [102]; using gravitational lensing data for DM mapping [103]; obtaining quasi-normal modes of black holes in General Relativity [104, 105, 106]; study of cosmological equations that govern the background dynamics of the Universe [107]; renormalization group flows on the lattice [108] and only very recently in the context of high-energy-physics [109, 110, 111, 112]. One type of generic problem PINNs have been used for are the so-called forward problems. Namely, given a physical law, usually characterized by PDEs or ODEs, and initial/boundary conditions, a solution is found by the PINN, corresponding to a physical quantity. But there is another interesting class of problems that can be solved by PINNs: inverse problems. In this case, the observable quantities are given to the PINN which subsequently determines the physical laws and/or parameters of a given physical theory that can satisfy such observables. This is of particular interest, since PINNs can potentially be a tool to discover models and/or a given models’ parameter space that inherently satisfy experimental observables.
@@
Inspired by the above ideas, in this work, we solve the paradigmatic freeze-in DM BE in alternative cosmologies using PINNs. Our goal is to showcase how PINNs are a powerful technology for exploring theoretical physics models motivated by experimental data in the context of particle DM – an approach that has so far not been explored in the literature. Specifically, we aim to address the following:
@@ 2 Freeze-in dark matter in alternative cosmologies
with $S_{0}=2.89\times 10^{3}$ cm-3 being the entropy at present time, $\rho_{\text{crit}}^{0}=1.05h^{2}\times 10^{-5}$ GeV///cm3 the present critical energy density and $h=H_{0}/(100$ km///s///Mpc) the reduced Hubble constant [113]. The above needs to be confronted with the experimentally observed value $\Omega_{\text{CDM}}h^{2}=0.120\pm 0.001$ [3].
@@ 3 Recasting the Boltzmann equation for freeze-in dark matter
The performance of the PINN can be further improved if we normalize $W(z)$ and $z$. Since the derivative $dY(x)/dx$ is always positive in Eq. (1) and the logarithm is a strictly monotonically increasing function, any initial condition (IC) we impose establishes the minimum of $W(z)$. Considering $z_{0}>0$ the point where such minimum occurs with $W_{0}\equiv W(z_{0})$, and denoting the final point as $z_{f}$, we rescale $W(z)$ and $z$ as follows,
$W_{n}(z)=-W(z)/W_{0}\;,\;z_{n}=z/z_{f}\;.$ (14)
The minimum of $W_{n}(z)$ is $-1$, while its maximum is within the interval [1,0[[-1,0[[ - 1 , 0 [.
@@ 4 Forward problem: parametric PINN solution
where we fix $W_{0}=\ln 10^{-25}$. In order to improve the performance of PINNs in solving PDEs, some mechanisms to make loss weights adaptive have been put forward, such as learning rate annealing [80, 115] and soft attention [116]. Although the BE imposes similar challenges to those tackled in the aforementioned references, we found no advantages in implementing these mechanisms. Hence, $\lambda_{\text{BE}}$ and $\lambda_{\text{IC}}$, were kept constant throughout training. Note that a full hyperparameter optimization is beyond the scope of this work, which aims to demonstrate the potential of PINNs as a valuable tool for particle DM studies. Thus, for our purposes, we select the weights and the learning rate via manual tuning. We found better results when $\lambda_{\text{IC}}$ was fixed at a higher value than $\lambda_{\text{BE}}$, so we take $\lambda_{\text{BE}}=1$ and $\lambda_{\text{IC}}=10$.
@@
Figure 3: Parametric forward PINN solution: particle yield evolution $W_{n}(z_{n})$ for Standard (top), RS (bottom left) and GB (bottom right) cosmology. Parametric dependence of $W_{n}(z_{n})$ on $C$ is shown via a colormap. The red-dashed contours represent FEM results for $C=-14$ and $C=-9$ in Standard cosmology across all panels. Horizontal dotted-black line corresponds to $W_{n}$ value satisfying $\Omega h^{2}=0.12$ [3].
Figure 4: Relation between $W_{n}(z_{nf})$ and $C$ for Standard (blue), GB (green) and RS (red) cosmology.
@@
In Fig. 3 we present the parametric PINN solution for the forward problem in Standard (top), RS (bottom left) and GB (bottom right) cosmology. The behavior of the yield $W_{n}(z_{n})$ with respect to $C$ is represented by a colormap, while the red-dashed lines show the solution provided by solve_ivp, when $C=-14$ and $C=-9$, for Standard cosmology. We observe a nearly perfect agreement between the PINN and solve_ivp in the top plot, with the solution for $C\in[-14,-9]$ filling practically all the region between the red-dashed lines. This showcases the accuracy of the forward PINN in solving the BE. Moreover, we notice that for $C\sim-11.5$, in Standard cosmology, our FIMP DM candidate accounts for the observed DM relic abundance (horizontal dotted-black line). For GB cosmology one would need a decreased value of $C\sim-13.5$ (bottom right plot), while for RS an increased value $C>-9$ is required (bottom left). These results are consistent with the theoretical expectations outlined in Section 2 [see Eqs. (9) and (12)]. In addition, in Fig. 4, we present the yield value provided by the PINN for the last collocation point $W_{n}(z_{nf})$ in terms of $C$, for the three cosmologies. The PINN confirms the expected linear relation between relic density and cross-section for FIMP DM [see Eqs. (9) and (12)]. Overall, compared to FEMs, the forward PINN provides a mesh-free method to solve the BE parametrically for cross-section values within an interval. Indeed, the NN is here a continuous function that approximates our target $W(z,C)$.
Figure 5: Forward PINN solution: evolution in terms of epochs of the mean relative absolute error between PINN and FEM $W(z)$ results [see Eq. (22)] (left) and total loss function $\mathcal{L}_{\text{Fwd}}$ [see Eqs. (19), (20) and (21)] (right), for Standard cosmology with $C=-11.6964$ (see Sec. 5.1).
To systematically benchmark the performance of the forward PINN against the FEM solver – which we take as our ground truth – it is useful to compute the mean relative absolute error:
$\bar{\varepsilon}_{\text{rel}}=\frac{1}{N_{z}}\sum_{j=1}^{N_{z}}\left|\frac{W(z_{j})^{\text{PINN}}-W(z_{j})^{\text{FEM}}}{W(z_{j})^{\text{FEM}}}\right|\;,$ (22)
between the PINN and FEM solution for the particle yield $W(z)$. This allows to impose an accuracy threshold on the PINN results defining a systematic criterion to assess convergence and stop training. As an example, we present in Fig. 5 the forward PINN solution for Standard cosmology with $C=-11.6964$ (see Sec. 5.1). Namely, on the left we plot the evolution in terms of epochs of the mean relative absolute error defined above, where it is clear that at the end of training our PINN reaches sub– $0.001\%$ agreement with the FEM solution. This result is consistent with the observed stabilization of the total PINN loss shown in the right plot, both plots exhibiting the same decreasing behavior in terms of the epochs. This is due to the fact that the mean relative absolute error is related to the residuals components of the loss function. These results showcase that in contrast to conventional NNs, where overfitting is a concern, PINNs typically improve with continued training.
Although we focus here on a benchmark with constant $\langle\sigma v\rangle$ to solve the freeze-in BE using PINNs, it is useful to outline possible future extensions towards more realistic scenarios. Namely, a natural avenue for future research would be to incorporate full Boltzmann collision integrals, which are essential for modeling realistic interactions involving multi-particle states and complex processes such as coannihilation, semi-annihilation, and inelastic scatterings. These arise in a wide range of well-motivated concrete particle DM models, for e.g. involving the scalar portal [10], requiring dedicated numerical tools such as the state-of-the-art packages MicrOmegas [64, 65] and DarkSUSY [66] to perform these computations. Thus, in order for PINNs to become a viable alternative to these widely used tools, future works should focus on incorporating the full Boltzmann collision integrals. As a first step with PINNs, it would be interesting to study one of the simplest realistic cases involving the thermally averaged cross section for $2\to 2$ annihilation processes [4]:
[…]
where $\sigma v(s)$ is the energy-dependent annihilation cross section which depends on the particle masses and interactions of a given model. One promising idea is to use a modular PINN approach [117] where we construct a dedicated network to learn the thermal integral shown above while a second network solves the BE for particle DM abundance. Namely, as an input the first network would receive $\sigma v(s)$ and learn how to model $\langle\sigma v\rangle(x)$. The latter would then enter the freeze-in BE which is solved by the second PINN. This separation into multiple PINNs could facilitate more targeted training and efficient inference of $\langle\sigma v\rangle(x)$ and $Y(x)$, potentially offering a scalable and flexible framework for exploring more realistic DM scenarios with full collision dynamics.
@@ 5.1 Discovering particle interaction cross sections
$\mathcal{L}_{\Omega h^{2}}=\left|W(z_{f})-W_{\Omega h^{2}}\right|\;,$ (25)
measures the absolute difference between the prediction of the PINN for the last collocation point $W(z_{f})$ and the observed DM relic density [see Eq. (13)]. Like $\lambda_{\text{BE}}$ and $\lambda_{\text{IC}}$, the weight $\lambda_{\Omega h^{2}}$ is an hyperparameter that should be tuned to optimize the PINN’s performance. As in the forward problem we have previously discussed, no improvements were found by making those loss weights adaptive. Manual adjustment leads us to set $\lambda_{\text{BE}}=1$ and $\lambda_{\text{IC}}=\lambda_{\Omega h^{2}}=10$.
@@
Figure 9: Inverse PINN solution: particle yield evolution $W_{n}(z_{n})$ for Standard (top), RS (bottom left) and GB (bottom right) cosmology. Black-dashed curves were obtained via FEM.
In Figs. 8 and 9, we present the PINN solution for the inverse problem of finding the FIMP DM particle yield evolution and determining the particle interaction strength for Standard (blue line), RS (red line) and GB (green line) cosmology. Namely, in Fig. 8, we show how $C$ evolves throughout training. The inverse PINN reaches the correct value for particle cross-sections after about $50-100$ thousand epochs for the three cosmologies being tested. Remarkably, Fig. 8 also shows that this can be achieved regardless of the initial $C$ being quite close (as in the case of the Standard cosmology) or far (as it happens in the alternative cosmologies) from the value that the PINN is supposed to discover.
For inverse PINNs, a direct FEM comparison is nontrivial, as FEMs are not inherently designed for inverse inference, requiring embedding into a computationally intensive parameter scan within a broader optimization or search framework. A full comprehensive comparison between PINNs and FEMs lies beyond the scope of this work and warrants a dedicated study. Nonetheless we wish to outline that a possible approach could involve constructing a fine mesh grid in parameter space [e.g., for $(z,C)$ or $(z,\gamma)$] and solve the BE at all the points of the grid to identify the parameter values that reproduce the data. However, this would transform the FEM into a brute-force parameter-scan algorithm, which lacks scalability and efficiency. Instead, a feasible benchmarking strategy that can be adopted is to reinsert the values of $C$ (or $\gamma$) obtained by the PINN into the FEM solver to compute the yield function $W(z,C)$ [or $W(z,\gamma)$]. This enables the evaluation of the mean relative absolute error, as defined in Eq. (22), in terms of the training epochs. In addition, the FEM-computed yield at the final collocation point, $W(z_{f})_{\text{FEM}}$, can be compared to the target value $W_{\Omega h^{2}}$ corresponding to the observed relic abundance, as a function of the training epochs. This strategy should provide a robust and reproducible quantitative metric for validating the inverse PINN solution. Instead, here we adopt a simple strategy to cross-check the PINN results which we now discuss. The PINN predictions for $C$ and corresponding cross-section [see Eq. (15)] are gathered in Table 1 for the three cosmologies. We also show the particle yield values for the last collocation point $W(z_{f})$ and relic density predictions stemming from our FEM BE solver when we provide it the $C$ values obtained by the PINN and their respective $\gamma$’s. This serves as a cross-check of the PINN results in obtaining the correct CDM relic abundance. In the last column of the table, the central value of the data is reproduced up to the fourth decimal place, showcasing the accuracy of the inverse PINN. Moreover, from Fig. 9, we also conclude that all cosmological scenarios reproduce the data $\Omega h^{2}=0.12$ (horizontal black-dashed line).
$\gamma$ $C_{\text{PINN}}$ $\langle\sigma v\rangle$ $\left(\text{GeV}^{-2}\right)$ $W(z_{f})_{\text{FEM}}$ $\Omega h^{2}_{\text{FEM}}$
@@
Table 1: Inverse PINN solution: finding $C$ [or equivalently $\langle\sigma v\rangle$ – see Eq. (15)] for GB ($\gamma=-2/3$), Standard ($\gamma=0$) and RS ($\gamma=2$) cosmology – see Figs. 8 and 9. Last two-columns indicate relic abundance values obtained via FEM.